# A Survey of Pun Generation: Datasets, Evaluations and Methodologies

### v1

@@ Abstract
Pun generation seeks to creatively modify linguistic elements in text to produce humour or evoke double meanings. It also aims to preserve coherence and contextual appropriateness, making it useful in creative writing and entertainment across various media and contexts. Although pun generation has received considerable attention in computational linguistics, there is currently no dedicated survey that systematically reviews this specific area. To bridge this gap, this paper provides a comprehensive review of pun generation datasets and methods across different stages, including conventional approaches, deep learning techniques, and pre-trained language models. Additionally, we summarise both automated and human evaluation metrics used to assess the quality of pun generation. Finally, we discuss the research challenges and propose promising directions for future work.
@@ 1 Introduction
{forest}
[…]
@@
The paper is organised as follows: Section 2 reviews the main categories of puns and provides examples for each category. Section 3, 4 and 5 summarise the relevant datasets, methods, and evaluation metrics, as shown in figure 1. We also discuss the challenges and outline future research directions in Section 6, as well as conclude with final remarks in Section 7.
@@ 5.1.2 Diversity
[…]
where a higher Dist-1 score indicates greater diversity in the generated sentences, whereas a lower score suggests more generic and repetitive text. The Dist-2 formulation is shown as Appendix B.
@@ Acknowledgments
This research is supported by the New Zealand Tertiary Education Commission and the Strong AI Lab at the University of Auckland. The first author of this research is funded by the China Scholarship Council (CSC).
@@ B.2 Supplement Details
$S_{ratio}\coloneqq\begin{cases}-1,&S_{local}<0\text{ or }S_{global}<0,\\ S_{local}/{S_{global}},&\text{otherwise}.\end{cases}$ (8)
@@
Dist-1 & Dist-2. We supplemented the formulation of Dist-2 here:
[…]
where a higher Dist-2 score indicates greater diversity in the generated sentences, whereas a lower score suggests more generic and repetitive text.

### v2

@@ Abstract
Pun generation seeks to creatively modify linguistic elements in text to produce humour or evoke double meanings. It also aims to preserve coherence and contextual appropriateness, making it useful in creative writing and entertainment across various media and contexts. Although pun generation has received considerable attention in computational linguistics, there is currently no dedicated survey that systematically reviews this specific area. To bridge this gap, this paper provides a comprehensive review of pun generation datasets and methods across different stages, including conventional approaches, deep learning techniques, and pre-trained language models. Additionally, we summarise both automated and human evaluation metrics used to assess the quality of pun generation. Finally, we discuss the research challenges and propose promising directions for future work.  [^1] https://github.com/ysu132/Pun-Generation-Survey
@@ 1 Introduction
Pun Generation Dataset Generic Datasets E.g. Wikipedia; BookCorpus Derived Datasets E.g. Hempelmann (2003) Bell et al. (2011) Human Annotated E.g. SemEval Miller et al. (2017) CUP Sun et al. (2022b) UNPIE Chung et al. (2024) Methods conventional Methods Manual E.g. Manurung et al. (2008) Tyler et al. (2020) Automatic E.g. Hong and Ong (2009) Valitutti et al. (2009) Classic DNNs E.g. Yu et al. (2018) Luo et al. (2019) Diao et al. (2020) Language Models Fine-Turning Auto-encoding E.g. Zeng et al. (2024) Auto-regressive E.g. Tian et al. (2022) Touvron et al. (2023) Encoder-decoder E.g. Mittal et al. (2022) Sun et al. (2022b) Prompting E.g. Xu et al. (2024a) Sarrof (2025) Visual-Language Models E.g. Rebrii et al. (2022) Chung et al. (2024) Evaluation Automatic Evaluation E.g. Dist-1 & Dist-2 Li et al. (2015) Ambiguity Kao et al. (2016) Human Evaluation E.g. Success Diao et al. (2020) Funniness Yu et al. (2020)
@@
Overall, the paper is organised as follows: Section 2 reviews the main categories of puns and provides examples for each category. Section 3, 4 and 5 summarise the relevant datasets, methods, and evaluation metrics, as shown in figure 1. We also discuss the challenges and outline future research directions in Section 6, as well as conclude with final remarks in Section 7.
@@ 5.1.2 Diversity
$\text{Dist-1}=\frac{unique\;unigrams}{total\;generated\;words}$ (5)
$\text{Dist-2}=\frac{unique\;bigrams}{total\;generated\;bigrams}$ (6)
where a higher Dist-1 and Dist-2 score indicates greater diversity in the generated sentences, whereas a lower score suggests more generic and repetitive text.
@@ Acknowledgments
This research is supported by the Strong AI Lab and the Natural, Artificial, and Organisation Intelligence Institute at the University of Auckland. The first author of this research is funded by the China Scholarship Council (CSC).
@@ B.2 Supplement Details
$S_{ratio}\coloneqq\begin{cases}-1,&S_{local}<0\text{ or }S_{global}<0,\\ S_{local}/{S_{global}},&\text{otherwise}.\end{cases}$